OpenStudy (anonymous):

HELP PLEASE, MEDAL TO BEST ANSWER!!! What is the value of the y variable in the solution to the following system of equations? 3x + y = 11 -3x + 2y = 1 Choose one answer. a. 2 b. 4 c. y can be any number as there are infinitely many solutions to the system. d. There is no y value as there is no solution to the system.

OpenStudy (anonymous):

@iGreen @AleshGames @AnswerMyQuestions @youngnephew34

Directrix (directrix):

@RedBlade127 3x + y = 11 -3x + 2y = 1 ========== Add the two equations. The x terms will add to zero and you can solve for y.

OpenStudy (anonymous):

Add them?

OpenStudy (igreen):

Yes..3x + -3x = 0 What's y + 2y? @RedBlade127

OpenStudy (anonymous):

Just y

OpenStudy (igreen):

No.. 'y' is the same thing as 1y. So what's 1y + 2y? (It's the same thing as 1 + 3 then adding a 'y' at the end of it)

OpenStudy (igreen):

Same thing as 1 + 2*

OpenStudy (anonymous):

3y

OpenStudy (igreen):

Yes, and what's 11 + 1?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

now what?

OpenStudy (igreen):

That gives us: 3y = 12 Now divide 12 / 3, what do you get?